OpenStudy (anonymous):

Fidelity Mutual (an insurance company) has offered you a single premium annuity that will pay you $12,000 at the end of each year for the next 15 years. If you must pay $109,296 today for this annuity, what is your expected rate of return?

OpenStudy (tkhunny):

Level payments? Why is that hard? You should know this. They only get harder. i = Expected Rate of Return v = 1/(1+i) \($109,296 = $12,000\cdot\dfrac{1 - v^{15}}{i}\) What's your plan for solving that fir 'i'?

OpenStudy (anonymous):

I came up with 9%, sorry, I didn't get all of my question entered. Is 9% correct?

OpenStudy (tkhunny):

How did you get 9%?

OpenStudy (anonymous):

divided both sides by 12,000

OpenStudy (tkhunny):

And then...

OpenStudy (anonymous):

divided both sides by i, then divided both sides by 1

OpenStudy (tkhunny):

Very bad. Let's follow the rules of algebra. 1) I hope you MULTIPLIED by 'i'. 2) There is no "divide both sides by 1". Why would that help, anyway? if 5x = 4, dividing both sides by 1 produces 5x = 4. Well, THAT didn't do anything!

OpenStudy (anonymous):

Sorry, I meant multiplied. I haven'e done this for 20 years and am trying to do work on another problem while we are working on this one. so now I have 9.108 (i) = 1/1+i correct

OpenStudy (anonymous):

9.108i

OpenStudy (tkhunny):

No. I'm still wondering how you are getting anywhere near 9%, since the correct value is 7%. I probably should let you off the hook. This equation CANNOT be solved explicitly for 'i'. It is an interative process. Have you EVER done that? Maybe you have a calculator that will do it for you?

OpenStudy (anonymous):

This is for a corporate finance class and interative process is not in any of the readings or examples. Thank you for your help, but I think I may have to ask the teacher for help which is also difficult because this is an online class.